shila is 15 years older than her brother sanjay. ten years from now shila will be twice as sanjay. find their present ages.
Answer by lwsshak3(11628)

shila is 15 years older than her brother sanjay. ten years from now shila will be twice as sanjay. find their present ages
**
let x=sanjay's present age
x+15=shila's present age
10 years hence:
sanjay's age=x+10
shila's age=x+15+10=x+25
x+25=2(x+10)=2x+20
x=25-20=5
x+15=20
ans:
sanjay's present age=5
shila's present age=20